OpenStudy (anonymous):

Find the remainder of the division problem shown below. f(x)=x^3+2x^2-6x-9 / x+2

OpenStudy (anonymous):

3\[\frac{x^3+2 x^2-6 x-9}{x+2}=\left(x^2-6\right)+\frac{3}{x+2} \]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Your are welcome.

OpenStudy (experimentx):

another method ... use remainder theorem \[ f(x) = x^3+2x^2-6x-9\] \[ f(-2) = (-2)^3 + 2(2)^2 - 6(-2) - 9 = 3 \] http://www.wolframalpha.com/input/?i=x^3%2B2x^2-6x-9+%2C+x%3D-2